### v3

# EC-LDA : Label Distribution Inference Attack against Federated Graph Learning with Embedding Compression

Abstract
Graph Neural Networks (GNNs) have been widely used for graph analysis. Federated Graph Learning (FGL) is an emerging learning framework to collaboratively train graph data from various clients. Although FGL allows client data to remain localized, a malicious server can still steal client private data information through uploaded gradient. In this paper, we for the first time propose label distribution attacks (LDAs [^1] The term “LDA” here is different from other machine learning terms like Latent Dirichlet Allocation.) on FGL that aim to infer the label distributions of the client-side data. Firstly, we observe that the effectiveness of LDA is closely related to the variance of node embeddings in GNNs. Next, we analyze the relation between them and propose a new attack named EC-LDA, which significantly improves the attack effectiveness by compressing node embeddings. Then, extensive experiments on node classification and link prediction tasks across six widely used graph datasets show that EC-LDA outperforms the SOTA LDAs. Specifically, EC-LDA can achieve the Cos-sim as high as 1.0 under almost all cases. Finally, we explore the robustness of EC-LDA under differential privacy protection and discuss the potential effective defense methods to EC-LDA. Our code is available at https://github.com/cheng-t/EC-LDA.
I Introduction
I-A Background
Graph Neural Networks (GNNs), designed to process graph-structured data, have gained significant attention for their effectiveness across various applications including recommendation systems [1], social networks [2], and protein interaction prediction [3]. GNNs can capture information between neighboring nodes, enhancing the expressiveness of node embeddings and making them highly suitable for real-world applications.
The performance of GNNs requires a large amount of data. However, due to privacy concerns and regulatory restrictions, machine learning platforms cannot access raw data directly, which makes centralized GNN learning challenging. In recent years, a lot of works [4, 5] have integrated Federated Learning (FL) [6] with GNNs, proposing the Federated Graph Learning (FGL). FL is a distributed, privacy-preserving machine learning paradigm that enables clients to train models collaboratively while keeping their local data isolated. It addresses the challenge of data silo, where data is distributed across different sources and cannot be easily combined for joint analysis, by enabling model training without the need to share the raw data.
However, recent studies have shown that local data in FL remains vulnerable to various attacks [7, 8, 9, 10]. One such attack is label distribution inference attacks (LDAs) [11, 12], which aim to infer the label distribution of client-side local training data by analyzing the gradients shared between clients and the server. This represents a significant privacy threat in FL. For example, if multiple online shopping companies collaborate to train a recommendation system model (as shown in Fig. 1), a malicious server with access to the label distribution of a private social network could target specific users, increasing the success rate of fraudulent activities and posing a serious threat to user privacy.
Figure 1: Threat model of EC-LDA.
I-B Previous Works and Limitations
Despite existing works on LDAs [11, 12, 13, 14], they all under non-GNN settings and have many limitations. For example, their attack methods require additional auxiliary datasets [11, 12] or specific activation functions [13] as assumption. Additionally, the performance of attack method declines as the local epochs in FL increase, achieving effective results only with a single local training epoch [14, 12]. These assumptions restrict their applicability to GNN scenarios. Furthermore, the message-passing characteristic among adjacent nodes in GNNs complicates node embeddings, potentially introducing irrelevant details into the model output, thereby making LDA in GNNs more challenging.
I-C Our Contributions
In this paper, we propose Embedding Compression-Label Distribution Inference Attack (EC-LDA), a novel LDA in FGL. Through exploring the factors behind the poor performance of LDAs in GNNs, we find there are strong correlation between the effectiveness of LDA and the variance of node embeddings. In GNNs, the variance in node embeddings primarily originates from the GNN layers and increases with the number of layers, which in turn degrades the attack performance of LDA. EC-LDA addresses this issue by compressing node embeddings, thereby reducing their variance and enhancing attack performance. Moreover, EC-LDA overcomes the drawbacks of existing LDAs. It keeps stable performance as training samples and local epochs increase and doesn’t need specific activation functions or extra datasets for attacker. Our main contributions are as follows:
  • We analyze the relationship between the performance of LDA and the message-passing characteristics of GNNs, and introduce EC-LDA, the first approach that implement efficient LDA on FGL.
  • We apply EC-LDA to six graph datasets and conduct extensive experiments targeting both node classification and link prediction tasks. These experiments demonstrate that EC-LDA consistently achieves significant attack performance across various scenarios. EC-LDA achieves a Cos-sim as high as 1.00 under almost all cases.
  • We utilize DP-GNN (node-level differential privacy) and Label-DP (label-level differential privacy) for local GNN training and evaluate the robustness of EC-LDA. We further discuss potential effective defense methods for EC-LDA.
I-D Paper Structure
The rest of the paper is organized as follows. Section II introduces the preliminaries knowledge and the threat model. In Section III, we analyze the message-passing in LDA and give the key observation. Though the findings, we proposed our attack EC-LDA approach in Section IV. The experimental results are presented in Section V. Section VI reviews related work, followed the limitations in Section VII. Lastly, conclusion in Section VIII.
II Preliminaries
II-A Graph Neural Networks
In general, GNNs are designed to process data $G(V,\mathbb{E})$ structured as a graph. Here, $V$ represents the set of nodes and $\mathbb{E}$ represents the adjacency matrix of $G$. Each node $v_{i}\in V$ has a feature vector $u_{i}$. GNNs generate valuable node embeddings via message-passing, making them suitable for various downstream tasks like node classification and link prediction. In this paper, we consider Graph Convolutional Networks (GCN) [15], Graph Attention Networks (GAT) [16], and GraphSAGE [17] as target GNNs.
GNNs typically follow the message-passing strategy that updates the features of nodes iteratively by aggregating the features of their neighbors. Typically, a GNN model’s $h$-th layer can be formulated as:
$u_{i}^{h}=\sigma(u_{i}^{h-1},AGG(u_{j}^{h-1},j\in{\mathbb{B}}_{i})),$ (1)
where $u_{i}^{h-1}$ is the representation obtained at the (h1)(h\!-\!1)-th layer of node $v_{i}$, and $u_{i}^{0}$ is the node feature $u_{i}$ of node $v_{i}$, ${\mathbb{B}}_{i}$ represents the neighbors of node $v_{i}$, $AGG(\cdot)$ represents the aggregation function, $\sigma$ represents the activation function such as $ReLU$.
TABLE I: ErrErr with the variance of II when local epochs $E$ is set to 1 (WikiCS dataset and a 2-layer GCN model).
Variance*1000 of II 41.052 23.719 21.283 10.640 4.547 2.031 0.817 0.224 0.067 0.003
ErrErr 4.367 3.433 2.765 2.217 1.523 1.012 0.599 0.309 0.138 0.028
II-B Federated Graph Learning
A typical FGL system follows the FedAvg [6] algorithm. Specifically, the server sends an initial global model to all the clients. Then each client trains a local model with its local private data $G_{i}(V_{i},\mathbb{E}_{i})$ and shares its local model parameters with the server. The server then aggregates the local model parameters of all the clients to construct the global model’s parameters, which can be formulated as:
$W^{t}=\sum_{i=1}^{N}p_{i}W_{i}^{t},$ (2)
where $N$ is the number of clients, $W_{i}^{t}$ and $p_{i}$ are trained model parameters at the $t$-th epoch and the weight of client $i$, respectively, $W^{t}$ is the global model parameters at the $t$-th epoch. The optimization problem of FGL is formulated as:
$W^{*}=\underset{W}{\operatorname*{argmin}}\sum_{i=1}^{N}p_{i}\mathcal{L}(W,V_{i},\mathbb{E}_{i};Y_{i}),$ (3)
where $\mathcal{L}(\cdot)$ represents the loss function, $Y_{i}$ is node label in the node classification task or link label in the link prediction task of client $i$, respectively, $V_{i}$ represents the set of nodes and $\mathbb{E}_{i}$ represents the adjacency matrix of $G_{i}$.
II-C Threat Model
As shown in Fig. 1, we consider the FGL scenario where a malicious server is curious about local private data and attempt to reconstruct the label distributions of the clients private data. For instance, consider multiple hospitals collaboratively training a GNN model to predict patients’ disease probabilities. Each hospital maintains local patient data graphs where nodes represent patients, edges correspond to medical relationships, and labels denote diagnosis results. By inferring hospital-specific label distributions, attackers could obtain sensitive regional health statistics, potentially revealing disease prevalence in specific healthcare jurisdictions.
Attacker’s Knowledge and Capability. In this scenario, the malicious server adversary can not only observe all gradient updates uploaded by clients but also actively tamper with global model parameters or the aggregation process [18, 19, 20].
III Message-Passing Issue in LDA
In this section, we will analyze how the message-passing mechanism in GNNs exacerbates LDA. We begin with an introduction to LDA. The prior analysis [21] has shown that, for the $k$-th sample in a batch: $\Delta W_{l}^{(FC)}/E\approx\frac{1}{K}\sum_{k}(p_{k,l}\sum_{m}I_{k,m}-y_{k,l}\sum_{m}I_{k,m})$, where $\Delta W_{l}^{(FC)}$ is the sum of the gradients of the $l$-th output unit in the last fully connected layer along the input dimension, $I_{k,m}$ is the input of the $k$-th sample at the $m$-th input unit of the last fully connected layer, $p_{k,l}$ is the post-softmax probability at index $l$ of the $k$-th sample, $E$ is the local epochs. Furthermore, we have: $\sum_{k}y_{k,l}\cdot\bar{I}\approx\sum_{k}p_{k,l}I_{k}-K\Delta W_{l}^{(FC)}/E$, where $I_{k}=\sum_{m}I_{k,m}$, and $\bar{I}$ is the mean value of $I_{k}$.
In order to separate out $\sum_{k}y_{k,l}$, we assume that $I_{k}$ is close to its mean value, i.e., $\bar{I}$. Based on this assumption, we have:
$d_{l}$ $=\sum_{k}y_{k,l}$ (4)
$\approx\frac{\sum_{k}p_{k,l}I_{k}}{\bar{I}}-\frac{K\Delta W_{l}^{(FC)}}{E\bar{I}},$
where $d_{l}$ is the number of nodes inferred with label $l(l=1,2,\ldots,L)$. With the number of each label, we can compute the label distribution, which we define as $D$. In FL, we assume that the server can get $E$ and $\Delta W_{l}^{(FC)}$ from clients. Additionally, the dummy training data randomly generated by the server can be used to estimate $K$, $p_{k,l}$, $I_{k}$, and $\bar{I}$. We named this method as LDA.
Figure 2: Illustrates the trend of variance of II with the number of GCN layers. The experimental results at this point are the normalized results.
Figure 3: Illustration of the EC-LDA.
Now we apply and analyze the LDA in GNNs. For the sake of description, we define $I:=[I_{1},I_{2},\cdot,I_{K}]$. Starting from the above assumption, i.e., $I_{k}$ is close to $\bar{I}$, we can find that the error arises mainly because we replace $I_{k}$ with $\bar{I}$, i.e., we consider that $\sum_{k}y_{k,l}I_{k}\approx\sum_{k}y_{k,l}\cdot\bar{I}$. We define the magnitude of the error as:
$Err$ $=\frac{1}{K}\sum_{l}\lvert\sum_{k}y_{k,l}(I_{k}-\bar{I})\rvert$ (5)
$=\frac{1}{K}\sum_{k}\lvert I_{k}-\bar{I}\rvert$.
We can intuitively see that if the variance of $I$ is smaller, the $Err$ is smaller and the performance of the attack is better, and vice versa. To verify our conjecture, we conduct experiments on FGL by performing attacks during mid-training round and analyze the variance of $I$ versus $Err$. Table I illustrates how $Err$ varies with the variance of $I$, demonstrating that the $Err$ declines as $I$’s variance declines, which confirms our suspicions.
As mentioned earlier, GNNs incorporate neighbor information through the message-passing mechanism. However, the message-passing mechanism also increases the variance of node embeddings. Moreover, the variance of the node embeddings increases with the number of GNN layers in the global model, and the node embeddings are directly related to $I$. We conduct experiments with GNN models featuring 1, 2, and 3 GCN layers and observe how the variance of $I$ changes with the number of GCN layers. Fig. 2 illustrates the variance of $I$ as a function of the number of GCN layers. We scaled the variance to [0,1] through min-max normalization:
$V_{new}=\frac{V-V_{min}}{V_{max}-V_{min}},$ (6)
which $V$ is the variance of $I$. This normalization does not change the trend that the variance in all datasets increases with the number of GNN layers. The empirical results indicate a positive correlation between the depth of GCN layers and the variance of $I$. Consequently, optimizing attack efficacy necessitates minimizing the variance of $I$.
Algorithm 1 Compression of node embeddings
Input: Global model parameters $G$, clipping threshold $C$
Output: The global model parameters $G^{\prime}$ after clipping and compression
[…]
2:for $p\in G$ do
3:  $\mathbb{N}=\mathbb{N}+\|p\|_{2}^{2}$
4:end for
[…]
6:for $p\in G,p^{\prime}\in G^{\prime}$ do
7:  $p^{\prime}\leftarrow p/max(1,\frac{\mathbb{N}}{C})$
8:end for
9:return $G^{\prime}$
Algorithm 2 EC-LDA against FGL (Server-side)
Input: Client number $N$, global rounds $R$, attack rounds AA
Output: Final global model $W_{R}$, attack result $\mathbb{L}$
1: Initialize global model $W_{0}$, attack result $\mathbb{L}$
2:for $r=1,2,3,\ldots,R$ do
3:  if $r\in A$ then
4:   $W_{r-1}^{{}^{\prime}}\leftarrow W_{r-1}$, clipping $W_{r-1}$ with Algorithm 1
5:   Generate dummy training data $D_{dummy}$
6:   Input $D_{dummy}$ into $W_{r-1}^{{}^{\prime}}$ to obtain $p_{k,l}$, $I_{k}$ and $\bar{I}$
7:  end if
8:  for $n=1,2,3,\ldots,N$ do
9:   Client $n$ performs local training and uploads $W_{r,n}$ and gradients
10:   if $r$ $\in$ AA then
11:    Calculate $D$ based on Equation 4
12:    Add $D$ to $\mathbb{L}$
13:   end if
14:  end for
15:  Server aggregates all local models $W_{r,n}$ to $W_{r}$
16:  if $r\in A$ then
17:   Replace $W_{r}$ with the initial model $W_{r-1}^{{}^{\prime}}$
18:  end if
19:end for
IV Our Methodology: EC-LDA
Based on the above analysis, we conclude that the key to reducing error lies in minimizing the variance of $I$, which can be effectively achieved by constraining the absolute values of $I_{k}$ by clipping the model parameters.
IV-1 Overview
We now provide the introduction to EC-LDA, an enhanced attack method targeting GNNs. Unlike LDA, EC-LDA makes full use of Equation 4 by clipping the parameters of the global model before distributing it to clients, which reduces the variance of node embeddings, thus making the variance of $I$ smaller.
As shown in Fig. 3, EC-LDA consists of three main steps:
  • Step 1: Clipping model. The server clips the initial model and sends it to clients to get the gradients. Algorithm 1 demonstrates the clipping method. First the $\ell_{2}$ norm $\mathbb{N}$ of the model is computed, and then $p/max(1,\frac{\mathbb{N}}{C})$, i.e., the clipped parameters, will replace the original parameters $p$. After receiving the global model sent by the server, the client trains the model using its own private training data and uploads the trained model to the server.
  • Step 2: Forward propagation of dummy data. The server first generates dummy training data, which is then fed into the initial model. Through forward propagation, the server obtains intermediate outputs $_{k}, I¯\bar{I}$ from the input of the final fully-connected layer. Simultaneously, the model’s output is processed by the softmax layer to derive $p_{k,l}$.
  • Step 3: Calculating distribution. Following local client training and forward propagation of dummy training data, the server obtains intermediate variables gradients, $p_{k,l}$, $I_{k}$, and $\bar{I}$. It then calculates the private label distribution of client-side data by applying Equation 4.
IV-2 EC-LDA
Algorithm 2 presents the server-side implementation of EC-LDA within the FGL framework. The boxed part represents the additional component of EC-LDA compared to the normal training process. First, the server initializes the model parameters and attack results (Line 1). If the server does not perform an attack in current round $r$, the training process is no different from the normal training process. Otherwise, the server saves the initial model $W_{r-1}$ for round $r$ as $W_{r-1}^{{}^{\prime}}$ and clip $W_{r-1}$ with Algorithm 1. The server generates dummy training data and feeds it into $W_{r-1}^{{}^{\prime}}$ to obtain $p_{k,l}$, $I_{k}$ and $\bar{I}$ (Line 4-6). Then the server sends the clipped model down to all clients, and each client trains on its local private data and uploads the trained model. After training, server calculates the label distribution $D$ with Equation 4, and saves $D$ into $\mathbb{L}$ (Line 11-12). The server uses the saved initial model $W_{r-1}^{{}^{\prime}}$ of round $r$ to replace $W_{r}$ (Line 17). This prevents propagation of clipping-affected models to subsequent rounds, thereby ensuring minimal performance impact from sporadic attacks. The training process of clients does not differ from the general training process, clients first receive the model $W_{r-1}$ from the server, then train the model $W_{r-1}$ with the local privacy training data for $E$ times and upload the trained model to the server. In other words, EC-LDA trades one round of training resources for one round of label distributions of client-side private training data.
IV-3 Extend to Link Prediction
Implementing EC-LDA requires GNN layers with a fully connected final layer and the use of the cross-entropy loss function, which makes EC-LDA particularly versatile within classification tasks on GNNs, such as node classification. Since the link prediction is a binary classification task, EC-LDA is still applicable to link prediction tasks as long as the last fully connected layer of the model has two output units and the cross-entropy loss function is used. For node classification, EC-LDA can attack the label proportions, while for link prediction, it can target the graph density.
V Experiments
In this section, we demonstrate the effectiveness of EC-LDA through answering the following three research questions:
  • RQ1 - How effective is EC-LDA under node classification and link prediction tasks with real-world graph datasets?
  • RQ2 - How robust is EC-LDA across different experimental parameters?
  • RQ3 - What is the performance of EC-LDA under differential privacy protection?
V-A Experimental Settings
Datasets: We conduct experiments on six widely used public datasets: Cora [22], CiteSeer [22], LastFM [23], Facebook [24], CoraFull [25], and WikiCS [26]. Cora, CiteSeer, WikiCS, and CoraFull are citation network datasets. Facebook and LastFM are social network datasets. We show the main features of these datasets in Table II. To demonstrate the effectiveness of EC-LDA under conditions with a large number of labels, we select the CoraFull dataset, which has up to 70 labels.
Hyper-parameter Settings: All datasets use the SGD optimizer. If not specified, all the experiments in this paper use the following setup: all at $E$ of 5, clipping threshold $C$ of 0.01, number of clients of 10, and attack on all the clients in the middle round. All the experiments in this paper are taken with different random seeds to repeat 5 times, and the results are averaged.
TABLE II: Statistics of datasets
Types Datasets #Nodes #Edges #Features #Classes
Citation
Network
CiteSeer 2120 7358 3703 6
Cora 2485 10138 1433 7
WikiCS 11311 297033 300 10
CoraFull 18800 125370 8710 70
Social
Network
Facebook 22470 342004 128 4
LastFM 7624 55612 128 18
TABLE III: Performance of EC-LDA and baselines.
Dataset GNN types Cos-sim JS-div
EC-LDA Infiltrator iLRG LLG* EC-LDA Infiltrator iLRG LLG*
CoraFull GCN 1.000 0.946 0.299 0.418 0.000 0.031 0.400 0.506
GAT 1.000 0.977 0.314 0.395 0.000 0.017 0.415 0.515
GraphSAGE 1.000 0.946 0.249 0.304 0.000 0.035 0.486 0.552
LastFM GCN 1.000 0.953 0.590 0.414 0.000 0.040 0.212 0.410
GAT 1.000 0.978 0.494 0.349 0.000 0.023 0.281 0.451
GraphSAGE 1.000 0.966 0.470 0.347 0.000 0.038 0.315 0.448
WikiCS GCN 1.000 0.865 0.610 0.475 0.001 0.058 0.212 0.402
GAT 1.000 0.987 0.622 0.488 0.001 0.010 0.214 0.396
GraphSAGE 1.000 0.721 0.561 0.512 0.001 0.130 0.242 0.374
Cora GCN 1.000 0.981 0.601 0.558 0.001 0.009 0.267 0.327
GAT 1.000 0.995 0.499 0.527 0.001 0.004 0.328 0.343
GraphSAGE 1.000 0.987 0.469 0.588 0.000 0.009 0.382 0.312
CiteSeer GCN 1.000 0.996 0.579 0.432 0.001 0.004 0.295 0.386
GAT 1.000 0.997 0.522 0.439 0.001 0.004 0.294 0.383
GraphSAGE 1.000 0.992 0.474 0.409 0.000 0.008 0.368 0.402
Facebook GCN 1.000 0.991 0.679 0.581 0.000 0.003 0.195 0.326
GAT 0.999 0.991 0.670 0.434 0.002 0.003 0.208 0.394
GraphSAGE 1.000 0.982 0.643 0.519 0.000 0.008 0.251 0.354
Evaluation Metrics: Inspired by [27], we use the following two evaluation metrics to fully demonstrate the effectiveness of EC-LDA: cosine similarity (cos-sim) and Jensen-Shannon divergence (JS-div). Cos-sim measures the similarity of two distributions and is applicable when the similarity of vectors is not directly related to the length of the vectors, while JS-div measures the distance between two distributions. The two metrics are calculated using the following two formulas:
$\text{JS-div}(D,D^{*})=\frac{1}{2}\sum_{n=1}^{N}d_{n}\log(\frac{2d_{n}}{d_{n}+d_{n}^{*}})$ (7)
$+\frac{1}{2}\sum_{n=1}^{N}d_{n}^{*}\log(\frac{2d_{n}^{*}}{d_{n}+d_{n}^{*}}),$
$\text{cos-sim}(D,D^{*})=\frac{\sum_{n=1}^{N}(d_{n}d_{n}^{*})}{\sqrt{\sum_{n=1}^{N}d_{n}^{2}}\sqrt{\sum_{n=1}^{N}d_{n}^{*2}}},$ (8)
where $D$ is the inferred label distribution and $D^{*}$ is the ground-truth label distribution. Cos-sim takes values in the range [1,1][-1,1]. The larger the cos-sim, the closer the inferred label distribution is to the ground-truth label distribution, i.e., the more effective the attack is. The value of JS-div ranges from $[0,1]$, and the more effective the attack is, the smaller the JS-div is.
Model Architecture: To illustrate the general applicability of EC-LDA, we choose three classical GNN models as global models, which are GCN, GAT, and GraphSAGE. All of the above models consist of two parts, the GNN layers and the fully connected layers.
Baselines: We compare EC-LDA with three different attacks which are Infiltrator [28], iLRG [14], and LLG* [12]. Infiltrator infers the label of the victim node by adding a neighbor to the victim node and observing the output of the neighbor. It is worth noting that Infiltrator focuses on node-level attacks, and for comparison, we attack all training nodes with Infiltrator. LLG* and iLRG both reveal the number of each label, and we use the number of samples per label extracted by iLRG and LLG* to compute the label distribution of each client. Similar to EC-LDA, iLRG and LLG* also use the gradients of the last fully connected layer of the model, and when $E$ is greater than 1, we use $\Delta W^{(FC)}/E$ as an approximation of the gradients.
Dummy Training Data: Based on the previous analysis, we know that the smaller the variance of $I$, the better the attack performance. Since the dummy training data is also related to the variance of $I$, we generate it from a Gaussian distribution with a mean of 0 and a standard deviation of 0.001, containing 1000 nodes.
V-B Attack Performance (RQ1)
Node Classification. We evaluate the performance of EC-LDA with all datasets and all GNN types, which, for the node classification task, have a varying number of labels and are widely distributed, with 4 kinds of labels for nodes in the Facebook dataset and 70 kinds of labels for nodes in the CoraFull dataset, the datasets with the smallest and the largest number of classes, respectively. Table III shows the experimental results. From the comparison presented in Table III, it is evident that EC-LDA consistently demonstrates exceptional performance across various datasets and three distinct GNN models, as indicated by the cos-sim scores consistently at or above 0.999 and the JS-div consistently at or below 0.002, aligning closely with the optimal values. Remarkably, EC-LDA’s outstanding performance remains consistent regardless of the number of labels, the number of nodes, and the specific GNN types, showcasing its broad applicability. Additionally, EC-LDA consistently outperforms other methods across all experiments.
TABLE IV: Performance of EC-LDA on link prediction tasks.
Dataset GNN types Cos-sim JS-div
Facebook GCN 1.000 0.000
GAT 1.000 0.000
GraphSAGE 1.000 0.000
CiteSeer GCN 1.000 0.000
GAT 1.000 0.000
GraphSAGE 1.000 0.000
Cora GCN 1.000 0.000
GAT 1.000 0.000
GraphSAGE 1.000 0.000
WikiCS GCN 1.000 0.000
GAT 1.000 0.000
GraphSAGE 1.000 0.000
LastFM GCN 1.000 0.000
GAT 1.000 0.000
GraphSAGE 1.000 0.000
CoraFull GCN 1.000 0.000
GAT 1.000 0.000
GraphSAGE 1.000 0.000
Link Prediction. In the link prediction experiments, we set the same number of positive and negative edges. Table IV shows the performance of EC-LDA on all datasets in the link prediction task. In all experiments of link prediction, cos-sim and JS-div reached 1.000 and 0.000, respectively, demonstrating the stunning performance of EC-LDA.
V-C Ablation Experiments (RQ2)
To explore the effectiveness of EC-LDA under varying parameters, we performed extensive ablation experiments on the node classification tasks. Specifically, we explored the impact of the GNN layer depth, local epochs $E$, clipping threshold $C$, number of clients, dummy training data separately. Furthermore, we also demonstrate the label distribution of clients during training.
Impact of the GNN Layer Depth. We illustrate in Fig. 4 the effect of GNN layer depth on the performance of EC-LDA. Specifically, we evaluate three GNN models with 1, 2, and 3 GNN layers, respectively. Fig. 4 shows that as the GNN layers increase, EC-LDA performance improves, contrary to Fig. 2. This is due to clipping reducing model parameters, which lowers the absolute values and variance of $I_{k}$, resulting in better attack performance. Overall, EC-LDA delivers excellent performance across three kinds of models.
Figure 4: Performance of EC-LDA under various the number of GNN layers.
Figure 5: Performance of EC-LDA under various clipping threshold $C$.
Figure 6: The label distribution obtained by each client when dividing the Cora dataset under five different random seeds.
Impact of clipping threshold $C$. In EC-LDA, the server will distributes the clipped model during attack rounds. The method of clipping directly affects the performance of EC-LDA. We employ the $\ell_{2}$ norm clipping method and investigate the impact of $C$ on EC-LDA. Fig. 5 shows that the performance of EC-LDA under different $C$. We can observe that as $C$ increases, the performance of EC-LDA deteriorates. This is because with increasing $C$, the clipping intensity decreases. When $C$ exceeds the $\ell_{2}$ norm of the model parameters, the value of $max(1,\frac{\mathbb{N}}{C})$ will be equal to 1, which means no clipping will be applied to the model, leading to a deterioration in EC-LDA’s performance.
Impact of local epochs $E$. In previous experiments we fixed $E$ to be 5, a widely accepted practice in FGL. To study the effect of different $E$ on EC-LDA, we evaluate the performance of EC-LDA under different $E$. We show the experimental results in Table V. It is evident that EC-LDA performs exceptionally well across varying $E$, with its performance across all datasets almost remaining unaffected by changes in $E$. As introduced in the previous section, we use $\Delta W_{n}^{(FC)}/E$ as an approximation of the gradients for EC-LDA. As $E$ increases, the error between the true gradients and the approximated gradients becomes larger, which results in a slight degradation of the performance of EC-LDA in some cases. Overall, EC-LDA performs effectively across different $E$.
Figure 7: Performance of EC-LDA under various client number.
Impact of the Client Number. To investigate the impact of the number of clients on the performance of EC-LDA, we conducted experiments on the Cora dataset, varying the number of clients from 5 to 50. Fig. 7 demonstrates the performance of EC-LDA as the number of clients changes. We can see that regardless of the number of clients, the performance of EC-LDA remains excellent, showcasing its wide applicability.
TABLE V: Performance of EC-LDA across various local epochs $E$.
Model
Local
Epochs
Cos-sim JS-div
Cora Facebook Cora Facebook
GCN 1 1.000 1.000 0.000 0.000
3 1.000 1.000 0.000 0.000
5 1.000 1.000 0.001 0.000
GAT 1 1.000 1.000 0.000 0.000
3 1.000 1.000 0.000 0.001
5 1.000 0.999 0.001 0.002
GraphSAGE 1 1.000 1.000 0.000 0.000
3 1.000 1.000 0.000 0.000
5 1.000 1.000 0.000 0.000
Impact of Dummy Training Data. In EC-LDA, dummy training data is employed to generate parameter $K$, $p_{k,l}$, $I_{k}$, and $\bar{I}$. Table VI illustrates the impact of the standard deviation in dummy training data on the performance of EC-LDA. The results indicate that changes in the distribution of dummy training data have minimal influence on the effectiveness of EC-LDA.
TABLE VI: Performance of EC-LDA under various standard deviation of dummy training data.
Dataset Standard Deviation Cos-sim JS-div
CiteSeer 0.001 1.000 0.001
0.1 1.000 0.001
10 1.000 0.001
Facebook 0.001 1.000 0.000
0.1 1.000 0.000
10 1.000 0.000
Impact of Label Distribution. We employ a community detection algorithm to partition the graph dataset for each client, which identifies communities within the network(i.e., groups of nodes with high connection densities). Fig. 6 shows the distribution of labels assigned to each client by this algorithm for the Cora dataset. From the figure, it is evident that, with a setup of 10 clients, the algorithm simulates a variety of label distributions. Our prior experimental results indicate that EC-LDA consistently delivers strong performance. Therefore, we conclude that EC-LDA can perform well across a range of distributions.
Figure 8: The attack performance of EC-LDA under no defense, DP-GNN defense, and Label-DP defense.
Figure 9: The model accuracy when executing EC-LDA attacks under no defense, DP-GNN defense, and Label-DP defense.
V-D Defense Performance of Differential Privacy (RQ3)
Differential privacy (DP) is a powerful privacy-preserving technology widely used in machine learning due to its rigorous mathematical definition. DP can defend against various attacks, such as membership inference attacks [29], adversarial example attacks [30], and data reconstruction attacks [31]. Specifically, we consider node-level differential privacy DP-GNN [32] and label-level differential privacy Label-DP [33]. DP-GNN controls the out-degree of nodes and is a variant of DP-SGD [34, 35]. Label-DP adds noise to the label matrix to protect the labels of the training dataset. In DP, privacy budget $\epsilon$ controls the scale of the added noise: a smaller privacy budget $\epsilon$ results in larger noise, and vice versa. Fig. 8 and Fig. 9 respectively present the performance of EC-LDA and the node classification accuracy of the FGL model under DP-GNN and Label-DP defenses.
Label-DP. From Fig. 8, it is evident that the defense effectiveness of Label-DP improves with decreasing privacy budget $\epsilon$. Under the Label-DP defense, noise is injected into the label matrix of the training dataset. As the $\epsilon$ value decreases, the magnitude of noise added to the label matrix progressively increases. This results in a widening divergence between the perturbed label matrix and the ground-truth label matrix. Additionally, the inferred label distribution closely matches the distribution after noise addition due to EC-LDA’s powerfulness in recovering label distributions. This alignment results in a significant disparity between the inferred distributions and actual label distributions, consequently diminishing EC-LDA’s attack effectiveness.
DP-GNN. However, as shown in Fig. 8, unlike Label-DP, the defensive efficacy of DP-GNN remains nearly unchanged as $\epsilon$ decreases from 5 to 1. Contrary to the well-established observation that diminishing $\epsilon$ enhances privacy protection, DP-GNN exhibits anomalous $\epsilon$-invariance. This unexpected phenomenon stems directly from the gradient processing mechanics of DP-GNN. As a variant of DP-SGD, DP-GNN first clips the gradients and then injects zero-mean Gaussian noise $\mathcal{N}(0,\sigma^{2})$ into gradients. However, when EC-LDA calculates the label distribution, the gradient it uses is $\Delta W_{l}^{(FC)}$, which is the sum of the gradients of the $l$-th output unit in the last fully connected layer along the input dimension, and $\Delta W_{l}^{(FC)}=\sum_{m}\Delta W_{m,l}^{(FC)}$. We define $\Delta\overline{W}_{l}^{(FC)}$ as the clipped version of gradient $\Delta W_{l}^{(FC)}$, and $\Delta W_{l}^{{}^{\prime}(FC)}$ as the gradient after adding noise to $\Delta\overline{W}_{l}^{(FC)}$. We have:
$\Delta W_{l}^{{}^{\prime}(FC)}$ $=\sum_{m}(\Delta\overline{W}_{m,l}^{(FC)}+\eta_{m,l}),$ (9)
where $\eta_{m,l}\sim\mathcal{N}(0,\sigma^{2})$, then $\mathbb{E}[\eta_{m,l}]=0$, in that way:
$\mathbb{E}[\Delta W_{l}^{{}^{\prime}(FC)}-\Delta\overline{W}_{l}^{(FC)}]$ (10)
$=\mathbb{E}[\sum_{m}\Delta\overline{W}_{m,l}^{(FC)}]+\mathbb{E}[\sum_{m}\eta_{m,l}]-\mathbb{E}[\sum_{m}\Delta\overline{W}_{m,l}^{(FC)}]=0.$
It can be interpreted as: during the summation of $\Delta W_{l}^{{}^{\prime}(FC)}$ over the input dimensions, the Gaussian noise added by DP-GNN undergoes positive-negative cancellation, causing the summed gradient after noise addition, i.e., $\Delta W_{l}^{{}^{\prime}(FC)}$, to approximate the clipped gradient, i.e., $\Delta\overline{W}_{l}^{(FC)}$. Consequently, this leads to the defensive efficacy of DP-GNN being insensitive to $\epsilon$.
VI Related Work
Inference Attack against GNNs. [36] investigated graph data leakage in horizontal federated and vertical federated scenarios. Specifically, link inference attack and attribute inference attack are proposed in the vertical federated scenario, graph reconstruction attack and graph feature attack are proposed in the horizontal federated scenario. [37] proposed an inference attack against the relationships between nodes of GNNs in vertical federated scenarios. [28] inferred the labels of victim nodes by infiltrating the raw graph data, i.e., by adding elaborately designed nodes and edges, and their attacks can only be performed on trained models. However, their attacks focus on the node-level, and there are no extant papers on label distribution inference attacks at the graph-level for FGL.
Label Recovery Attack. The label distribution inference attack is a type of label recovery attack. [7] was the first to restore the training sample from gradients. They restored the input data and associated label from the gradients using gradient-matching. This method continuously optimizes the dummy input data and associated label by minimizing the mean square error of the gradients of the dummy sample with respect to the true gradients. [38] introduced iDLG and were the first to propose that with a non-negative activation function, privacy label can be extracted with 100% success rate from the signs of gradients at the output layer. Both of these methods are applicable only to single-sample training scenarios. Afterwards, [13, 21] extended the attacks to the mini-batch scenario. Extensive research such as [14, 39, 40, 12, 11] demonstrates that gradients can be exploited to extract private label distributions. [41] relies on number theory and combinatorics to recover label information from log-loss scores. However, their method is sensitive to the number of classes in the dataset, and as the variety of labels increases, not only does the accuracy decrease, but the inference time also increases. [42] proposed three attack methods, which can infer whether a specific label appears in the training process, the quantity of each label for a specific client in a round, and the quantity of each label throughout the entire training process, respectively.
Overall, most existing label recovery attacks focus on image datasets, and none of them discuss label distribution on graph datasets. In this paper, We focus on label distribution inference attack against GNNs.
VII Limitations
In EC-LDA, the malicious server enhances attack effectiveness by clipping to compress the variance of node embeddings, then distributing this clipped model parameters to clients. This clipping operation leads to an anomalous distributional compression of model parameters. Consequently, a potentially effective defense strategy is for clients to detect potential attacks by identifying these abnormally compressed global model parameters and implementing corresponding countermeasures. However, since EC-LDA only requires a single clipping operation and can be executed at any point during the FGL process, this presents significant challenges for client-side detection and defense. For future work, we will explore effective detection and defense methods against EC-LDA. Additionally, we will concentrate on enhancing the stealthiness of our attacks. To enhance stealth, the server can deploy a “fishing model” that mimics the original model’s performance and parameter distribution while achieving the same effect as embedding compression, allowing the attack to proceed covertly while maintaining its effectiveness.
VIII Conclusion
This paper presents EC-LDA, a label distribution inference attack designed for Federated Graph Learning (FGL). The proposed method was developed by investigating the impact of node embedding variance in Graph Neural Networks (GNNs) on existing LDA approaches. By compressing node embeddings, EC-LDA significantly enhances the attack capability of a malicious server. Extensive experiments on six graph datasets show that EC-LDA outperforms state-of-the-art methods in both node classification and link prediction tasks. Additionally, we demonstrate the robustness of EC-LDA under various parameter settings and under DP protection.
IX ACKNOWLEDGMENTS
This work is supported by Fundamental Research Funds for the Central Universities (Grant No. 40500-20104-222609), National Natural Science Foundation of China Key Program (Grant No. 62132005), Natural Science Foundation of Shanghai (Grant No.22ZR1419100), and CAAI-Huawei MindSpore Open Fund (Grant No.CAAIXSJLJJ-2022-005A).